OpenStudy (anonymous):

How do you turn this into a degree (with decimals)? 55 degree 30' 10"????

OpenStudy (perl):

hint: there are 60 minutes in a degree , and 60 seconds in a minute

OpenStudy (perl):

$$ \Large 55^o~ 30'~ 10'' = 55 + \frac{30}{60} + \frac{10}{60^2} $$